OpenStudy (anonymous):

A recursive function is shown below: f(1) = 4 and f(n) = f(n - 1) - 8; n > 1 Which of the following lists the terms in the sequence defined by this recursive function?

OpenStudy (anonymous):

4, -4, -12, -20, -28, ... 4, 4, 12, 20, 28, ... 4, 12, 20, 28, 36, ... 4, -12, -20, -28, -36, ...

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

@mathmale @undeadknight26

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

I'm leaning toward A

OpenStudy (superhelp101):

steps 1. f(2) = f(2-1)-4 = f(1) - 4 = ? 2. So since you have f(1). 3. f(1) - 4 4. f(1) = 3. 5. f(1) - 4 = 3-4 = -1.

OpenStudy (anonymous):

What do you think the answer is?

OpenStudy (superhelp101):

I agree with you on A

OpenStudy (anonymous):

Okay, thanks!